OpenStudy (anonymous):

Find the 50th term of the sequence 5,-2,-9,-16

OpenStudy (***[isuru]***):

do u what type of sequence is this ? (arithmetic or geometric)

OpenStudy (anonymous):

arithmetic

OpenStudy (anonymous):

50+(49)*(-7) is the 50th term

OpenStudy (***[isuru]***):

now... can u find the common difference then ?

OpenStudy (anonymous):

yes, i got -293 is that right?

OpenStudy (anonymous):

wait... where did you get the 49 and the -7 from?

OpenStudy (***[isuru]***):

common difference(d) = difference between 2 successive terms = -2 -5 = -7 now \[A_{n} = a + ( n-1)d\] where simple "a" = first term = 5 n = the position of the term u want to find = 50 th term d = (-7) plug them in to the equation

OpenStudy (anonymous):

okay give me a sec

OpenStudy (***[isuru]***):

\[A_{50} = 5 + [ ( 50 - 1) \times (-7) ]\] \[A_{50} = 5 + ( 49 \times (-7) )\]

OpenStudy (anonymous):

i got -338

OpenStudy (anonymous):

can you help me with another one?

OpenStudy (***[isuru]***):

yep..! its the 50th term

OpenStudy (***[isuru]***):

sure.. if i can :)

OpenStudy (anonymous):

okay how can i find the second and third term in this sequence -7,___,___,-22,-27

OpenStudy (***[isuru]***):

subtract -22 from -27 -27 - (-22) = -27 + 22 = -5 its the common difference now add the common difference to the first term it will give u the second term ... now add the common difference to the second term and it will give u the third term .. can u do that ?

OpenStudy (anonymous):

i was so frustrated, this is easy lol